# Just-in-time and distributed task representations in language models

### v1

@@ 3.2 Task representations exhibit temporal locality
The analyses above confirm that, intuitively, in-context task representations can successfully benefit from increasing evidence in context (along the “temporal” dimension), and converge to better representations that restore task contexts more robustly. To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
@@
We extracted task vectors from the format tokens in the prompt, including "Q", the ":" following "Q", "A", and the new-line token before the "A". We tested whether these extracted representations can also restore the corresponding task contexts when patched onto the forward pass in the zero-shot prompt (as before, we patched at the same layer, but onto the corresponding format token instead of the last colon). As shown in Figure 2B, transferrable task representations generally do not robustly form in the residual activations across layers in these tokens, even though these tokens are also shared across examples, tasks, and contexts. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of in-context task representations shown in Figure 1C.
We observed nearly zero recontextualized zero-shot accuracy for all these tokens in most tasks, except some restoration success in product-company, color_v_animal_3, choose_first_of_5, and the longer-generation tasks discussed below (see Figure S3). In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token just before answer generation for each query.
@@
Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across the tasks and contexts. To study this, we trained simple, single-layer linear decoders to predict the task category (among the eleven simple tasks that we observed evidence accrual) from the residual activations of each format token. We report the test accuracy on the 25% held-out token activations extracted following eight examples in Figure 2B and Appendix C, but found high task identity decoding accuracy across these tokens following different numbers of examples. This decodability success may be partly due to vocabulary differences for the tasks, but the accuracy is close to ceiling, and not all tasks are distinguishable in this way. We note that task identity is not perfectly decodable in the token activations in early layers, but generally reaches high accuracy in mid and late layers. This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
@@ 3.3 Task representations exhibit scope locality
Figure 4: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over longer generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Left: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks. Right: recontextualized accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation – that is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? Prior work has mostly focused on simple, single-token output tasks. To study the semantic scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
@@
In these experiments, we find further evidence of the locality of in-context task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 4, left; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 4, right), suggesting that the restored task contexts “fade” over longer generation.
Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”. For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens. This effect is very pronounced in the mixed-generation tasks, as all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain and defer representing later subtasks. Interestingly, the extracted task representations that support generating responses for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 3B). This can potentially reflect the inert sensitivity to the broader mixed-task context, as well as conveying information relevant for later subtasks to be activated, even though these signals are not sufficient to directly restore execution of later subtasks.
The restored task contexts do support longer generation to some extent in some tasks, including repeating a simple task multiple times and the list-level tasks (i.e. reversing or shifting a word list). Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
@@ 4 Related work
From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025). More recently, attention has turned to cases where models may create internal task representations, that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt. Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information. Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024). Other works have explored how these representations and structures emerge over training (Yang et al.,, 2025; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024). Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
@@ 5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task context in zero-shot settings. Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts. However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask). The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context. In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
@@
In general, our results complicate the intuitive picture that the computational process underlying language model’s in-context evidence accrual and task inference is smooth and gradual. Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens. That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information at multiple layers may suggest a general inductive bias to compress knowledge to local representations when possible. This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
@@
However, one trend that arose from these investigations is that language models also do not seem able to condense task information into local representations in all cases. Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time. As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025). In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass. It is also possible that by overriding token activations with task vectors in intermediate layers, the models may have lost any task state information formed in earlier layers, and other methods that restore task states through additive injection rather than activation patching may be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024). Understanding how more complicated in-context computation develops based on evidence in these tasks and whether vector-form task representations can re-activate these inferences are interesting open questions.
@@
An intriguing direction for future work is to study whether there are mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. Yang et al.,, 2025).
@@ Appendix A Tasks
Appendix B Additional figures
Figure S1: Recontextualization accuracy for all tasks. Task vectors extracted from 8-shot prompts are used to reinstantiate task contexts in zero-shot settings. The dotted line indicate the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
Figure S2: Developmental trajectory of task representations over shots. Task representations are the token activations of the colon token prior to answer generation. We visualize task vectors sourced from the mode best layer across tasks at which task contexts are best restored in a zero-shot setting. Representations are first averaged across samples with the same number of examples for each task.
Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different models (see Figure S1).
Figure S4: Task representations across all models and tasks.
Appendix C Additional results
Table 4: Decoding accuracy for task identity at different tokens. Results shown for token activations extracted at the 8th example.
[…]
model layer
GEMMA_V3_4B_PT 2 0.933 0.848 0.925 0.899 0.950
5 0.979 0.980 0.816 0.925 0.994
8 0.952 0.986 0.908 0.959 0.996
11 0.950 0.986 0.968 0.984 0.998
14 0.960 0.990 0.988 0.984 0.996
17 0.902 0.994 0.990 0.996 0.996
20 0.918 0.992 0.970 0.982 0.992
23 0.834 0.988 0.992 0.954 0.996
26 0.842 0.865 0.990 0.916 0.992
29 0.811 0.973 0.992 0.922 0.986
32 0.905 0.948 0.992 0.958 0.966
GEMMA_V3_12B_PT 2 0.945 0.896 0.858 0.808 0.963
5 0.886 0.965 0.872 0.959 0.905
8 0.975 0.894 0.866 0.889 0.979
11 0.958 0.940 0.980 0.998 0.986
14 0.980 0.960 0.980 0.988 0.998
17 0.979 0.986 0.956 0.946 0.998
20 0.977 0.990 0.968 0.942 0.998
23 0.963 0.998 0.994 0.933 0.984
26 0.907 1.000 0.998 0.848 0.992
29 0.889 0.994 0.982 0.859 1.000
32 0.924 0.962 0.984 0.742 1.000
35 0.845 0.956 0.952 0.869 0.994
38 0.773 0.942 0.980 0.808 0.986
41 0.756 0.869 0.920 0.772 0.990
44 0.817 0.928 0.984 0.953 0.990
47 0.956 0.971 0.970 0.928 0.968
GEMMA_V3_27B_PT 5 0.929 0.876 0.829 0.879 0.905
11 0.940 0.994 0.865 0.946 0.972
17 0.979 0.990 0.900 0.996 0.996
23 0.973 1.000 0.945 0.960 0.996
29 0.950 0.998 0.994 0.937 0.994
35 0.864 0.988 0.924 0.906 0.994
41 0.936 0.964 0.960 0.827 0.990
47 0.890 0.996 0.912 0.858 0.990
53 0.887 0.980 0.912 0.787 0.968
59 0.863 0.984 0.994 0.977 0.958

### v2

@@ 3.2 Task representations exhibit temporal locality
The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations. To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
@@
We tested whether extracted representations from other format tokens in the prompt can also restore the corresponding task contexts. These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts. As before, we patched the activations at the same layer, but onto the corresponding format token instead of the last colon token in zero-shot prompts. As shown in Figure 3B, transferrable task representations generally do not form in the residual activations across layers in these tokens. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of transferrable task representations shown in Figure 1C.
We observed nearly zero recontextualized zero-shot accuracy for all these tokens in most tasks, except some restoration success in product-company, color_v_animal_3, choose_first_of_5, and the longer-generation tasks discussed below (see Figure S3). In Qwen3 models, there also appear to be some restoration successes with representations extracted from the “Q” token (see Figure S7B and Figure S11). These partial successes are likely driven by the fact that the first answer token sometimes match the first input token across a subset of the tasks. In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token before answer generation.
@@
Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across all tasks. We report the decoder accuracy from the layer with the best transferrable task representations in Figure 3B, but found high task identity decoding accuracy across most tokens throughout the context (Figure 2B). The decodability success may be partly due to vocabulary differences between some tasks, but we show that decoding accuracy remains high even for a restricted subset of tasks with shared vocabulary (Figure S4C). Both task identity and task transferability do not occur until at least one full example is presented, but accurate task identity representations form much earlier than transferrable task representations (Figure 1C). This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
@@ 3.3 Task representations exhibit scope locality
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation. That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? Prior work has mostly focused on simple, single-token output tasks. To study the scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
Figure 5: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks. Line plots: recontextualization accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
@@
In these experiments, we find further evidence of the locality of transferrable task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts “fade” over longer generation.
Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”. For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens. This effect is very pronounced in the mixed-generation tasks. In these cases, all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain, and defer representing later subtasks. Interestingly, the extracted task representations for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 4B), even though they support generating the same responses. This can potentially reflect an inert sensitivity to the broader mixed-task context based on vocabulary differences. Additionally, these differences may also convey information relevant for later subtasks to be activated, even though they are not sufficient to directly restore execution of later subtasks.
The restored task contexts do support longer generation to some extent in many tasks, including repeating a simple task multiple times and some list-level tasks (i.e., reversing or shifting a word list). Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
@@ 4 Related work
From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025). More recently, attention has turned to cases where models may create internal task representations. These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt. Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information. Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024). Other works have explored how these representations and structures emerge over training (Yang et al., 2025b, ; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024). Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
@@ 5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task contexts in zero-shot settings. Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts. However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask). The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context. In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
@@
In general, our results complicate the intuitive picture that the computational process underlying language models’ in-context learning and task inference is smooth and gradual. Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens. That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information may suggest a general inductive bias to compress knowledge to local representations when possible. This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
@@
However, one trend that arose from these investigations is that language models do not seem able to condense task information into local representations in all cases. Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time. As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025). In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass. It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost any task state information formed in earlier layers. Other methods that restore task states through additive injection rather than activation patching may therefore be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024). We also note that different models may differ in the extent to which they form local or distributed effective representations across the context or layers.
@@
An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. Yang et al., 2025b, ).
@@ Appendix A Tasks
Appendix B Additional Gemma3 results
Figure S1: Recontextualization accuracy for all tasks. Task vectors extracted from 8-shot prompts are used to reinstantiate task contexts in zero-shot settings. The dotted line indicates the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
Figure S2: Developmental trajectory of task representations over shots. Task vectors are the token activations of the colon token prior to answer generation. We visualize task vectors sourced from the mode best layer across tasks at which task contexts are best restored in a zero-shot setting. Representations for each task are first averaged across samples with the same number of examples in the prompt.
Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt in different tasks. The colors indicate different model sizes: blue=GEMMA_V3_4B_PT, yellow=GEMMA_V3_12B_PT, red=GEMMA_V3_27B_PT.
Figure S4: The development of different task representations over layers and context. A. Recontextualized accuracy when each token representation is used to restore task context in zero-shot settings. B. Task identity decoding accuracy across all 14 simple/short-answer tasks in Table 1. C. Task identity decoding accuracy across 9 list operation tasks with shared vocabulary. This subset of tasks includes: choose_first_of_5, choose_middle_of_5, choose_last_of_5, choose_first_last_of_3, choose_first_middle_last_of_5, reverse_all_of_3, reverse_all_of_5, shift_all_of_3, shift_all_of_5.
Figure S5: T-SNE projection of all task vectors across models and tasks.
Appendix C Results on Qwen3 models
Figure S6: Recontextualization accuracy from Qwen3 models for all tasks. Visualization as in Figure S1.
Figure S7: In Qwen3 models, transferrable task representations also condense evidence and activate sporadically at specific tokens. Visualization as in Figures 3A, B and Figure 1C. The hard-to-transfer tasks in Qwen3 models include english-french, choose_first_of_5, count_color_v_animal_3, and count_fruit_v_animal_3. These are the tasks where task vectors extracted from 32-shot prompts failed to recover more than 50% of the 32-shot accuracy on 0-shot prompts, across all model sizes.
Figure S8: Variance and magnitude changes across task vectors extracted following different number of examples. Visualization as in Figure 4A.
Figure S9: T-SNE task representations across all models and tasks.
Figure S10: Developmental trajectory of task representations over shots. Visualization as in Figure S2.
Figure S11: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different model sizes: blue=QWEN3_4B, yellow=QWEN3_8B, red=QWEN3_14B.
Figure S12: Task identity decoding accroacy across all 14 simple tasks in Table 1 (A) and 9 list operation tasks with shared vocabulary (B). See Figure S4 caption for a complete list of tasks.
Figure S13: In Qwen3, restored task contexts in longer- and mixed-generation tasks also decay over more output units during generation. Visualization as in Figure 5.